OpenStudy (anonymous):

What is the solution to 3 times the square root of 4 minus 2 times the square root of 4? A.0 B.1 C.2 D.3

OpenStudy (ddcamp):

Remember, sqrt(4) = 2: \[3\sqrt{4}-2\sqrt{4} = 3(2) - 2(2)\]

OpenStudy (anonymous):

Thank you!